Use partial fraction to evaluate inverse Laplace Transform of S- 1 F(s) = (s + 1)(s + 2)
Answer #1

Given function , f(3) = (5+ 1) (8 +2) - S-1 (S+1) 6+2) A (5+1) B (5+2) : A(S+2) + B(5+1) (S +1) C+2) = (A+B)s +(2A+B) $+1) (S